# Bench2Drive: Towards Multi-Ability Benchmarking of Closed-Loop End-To-End Autonomous Driving

### v2

@@ 5 Conclusion
Limitations: Bench2Drive is only in the simulation - CARLA, which could have gaps with real world. Future works might combine generative models to conduct simulation with authentic rendering.
Social Impact: Bench2Drive could serve as a platform to validate the skills of a AD system. Before putting into production, more rigorous road tests should be conducted to avoid safety issues.
@@ Appendix A Details of Data Collecting
Appendix B Distribution of Towns and Weathers
@@ Appendix C Implementation Details of Baselines
Appendix D Description of Scenarios
@@ 44. TJunction
Appendix E Details about Infraction Score
Infraction Penalty Note
Pedestrian Collision 0.50 Punished every infraction.
Vehicles Collision 0.60 Punished every infraction.
Other Collision 0.65 Punished every infraction.
Running Red Light 0.70 Punished every infraction.
Scenario Timeout 0.70 Fail to pass certain scenarios in 4 minutes.
No Give Way 0.70 Failure to yield to emergency vehicle.
Off-road - Not considered in route completion.
Agent Blocked - No action for 180 seconds. Shutdown immediately.
Route Timeout - Exceed the maximum time limit. Shutdown immediately.
Appendix F Author Statement
@@ Appendix G License
Appendix H Datasheet
H.1 Motivation
@@ H.4 Composition
How many instances are there in total (of each type, if appropriate)? There are 10000 clips in Bench2Drive.

### v3

@@ 5 Conclusion
Limitations: Since the rendering of simulation in CARLA has gaps compared to real world, utilizing real world datasets could be complementary as done in the concurrent work - NAVSIM [65]. Actually, there is a dilemma in the field for the evaluation of end-to-end autonomous driving algorithms:
Source of Images Pros Cons
Real World Datasets Realistic Non-Reactive
Simulation Rendering Reactive Cartoon Style
Generative models like diffusion models [66] might have the potential to provide realistic and reactive rendering, with some pioneering works in the field [67, 68, 69]. However, the illusion and artifact issue of diffusion requires further exploration.
Social Impact: The deployment of AD systems holds immense potential to revolutionize transportation, but it also brings significant ethical and safety concerns. Bench2Drive could serve as a platform for rigorously validating the capabilities of AD systems in a controlled and simulated environment, helping to identify potential flaws before real-world deployment. One of the primary risks is the simulation-reality gap—the difference between how an AD system performs in simulation versus in the real world. Simulations have the difficulties to fully replicate the complexities and unpredictability of real-world driving conditions. There is a risk that an AD system might perform well in simulation but fail in real-world scenarios due to unmodeled factors like rare edge cases, unexpected human behaviors, or varying environmental conditions. Bench2Drive is intended to complement, not replace, real-world testing, and it is crucial to emphasize that simulation is one part of a broader validation process that must include extensive on-road testing.
@@ Appendix A Details of Data Collecting
Appendix B Distribution of Scenarios, Towns and Weathers
@@
Figure 6: Scenario Distribution of Bench2Drive Dataset
@@ Appendix C Implementation Details of Baselines
Appendix D Training and Evaluation Resource Requirements
We report resource requirements of training and evaluating baselines. Note that the evaluation time could be linearly speed up with more GPUs to parallely evaluate on more routes.
Table 5: Resource requirements for training on base set and evaluation on 220 routes. The training of UniAD consists of BEVFormer, stage1, and stage2.
Training Evaluation
AD-MLP 1 A600 * 1 day 4 A6000 * 6 hours
TCP 1 A600 * 1 day 4 A6000 * 6 hours
VAD 8 H800 * 5 days 8 H800 * 2 days
UniAD-base 8 H800 * (3+3+3)=9 days 8 H800 * 2 days
Appendix E Behavior Model of NPC Agents
In CARLA, three behavior types are preset in CARLA Agent: cautious, normal and aggressive. These behavior types govern the driving actions of NPCs, influencing factors such as speed, responses to other vehicles, and safety protocols. The key parameters for each behavior mode include:
  • max_speed: Sets the maximum speed(km/h) that an NPC vehicle can reach.
  • speed_lim_dist: Value in km/h that defines how far your vehicle’s target speed will be from the current speed limit
  • speed_decrease: Controls the deceleration of the NPC when approaching a slower vehicle ahead.
  • safety_time: Estimates the time to a collision if the vehicle in front suddenly brakes.
  • min_proximity_threshold: Defines the minimum distance before the NPC takes actions such as evasive maneuvers or tailgating.
  • braking_distance: The distance at which the NPC performs an emergency stop to avoid a collision.
  • tailgate_counter: A counter that prevents the NPC from initiating a new tailgating action too soon after the last one.
These behavior designs interact with the ego (self-driving) vehicle, ensuring that NPCs respond to the presence and actions of the ego vehicle in the simulation. The parameters for different behavior styles are as follows,
Table 6: Comparison of Cautious, Normal, and Aggressive behavior parameters.
Parameter Cautious Normal Aggressive
max_speed (km/h) 40 50 70
speed_lim_dist (km/h) 6 3 1
speed_decrease (km/h) 12 10 8
safety_time (seconds) 3 3 3
min_proximity_threshold (meters) 12 10 8
braking_distance (meters) 6 5 4
tailgate_counter (times) 0 0 -1
The rule-based behavior decision algorithm in behavior_agent.py for non-player character (NPC) vehicles is shown below:
Algorithm 1 Non-Player Character (NPC) Vehicles Behavior Decision
1:  Update vehicle’s surrounding information
2:  if red light or stop sign detected then
3:     return  emergency_stop()
4:  end if
5:  if pedestrian detected and within braking distance then
6:     return  emergency_stop()
7:  end if
8:  if other vehicle nearby then
9:     if within braking distance then
10:        return  emergency_stop()
11:     else
12:        return  car_following()
13:     end if
14:  else if ego vehicle at intersection and (turn left or turn right) then
15:     adjust_speed_limit()
16:     return  PID_Control
17:  else if in normal driving conditions then
18:     maintain_speed_limit()
19:     return  PID_Control
20:  end if
The behavior of walkers/pedestrians is characterized by their consistent adherence to the route line at a constant speed, with the inability to walk backward. In autonomous driving scenarios, pedestrian safety is of utmost importance. Pedestrians have the highest priority when autonomous vehicles interact with humans, and autonomous vehicles should learn to give way to pedestrians regardless of traffic conditions.
Appendix F Details about Infraction Score
Infraction Penalty Note
Pedestrian Collision 0.50 Punished every infraction.
Vehicles Collision 0.60 Punished every infraction.
Other Collision 0.65 Punished every infraction.
Running Red Light 0.70 Punished every infraction.
Scenario Timeout 0.70 Fail to pass certain scenarios in 4 minutes.
No Give Way 0.70 Failure to yield to emergency vehicle.
Off-road - Not considered in route completion.
Agent Blocked - No action for 180 seconds. Shutdown immediately.
Route Timeout - Exceed the maximum time limit. Shutdown immediately.
Appendix G Description of Scenarios
@@ 44. TJunction
Appendix H Author Statement
@@ Appendix I License
Appendix J Datasheet
J.1 Motivation
@@
How many instances are there in total (of each type, if appropriate)? There are 13638 clips in Bench2Drive.